david currently has a square garden. He wants to redesign his garden and make it into a rectangle with a length that is 3 feet shorter than twice its width. He says that the perimeter is sixty feet. Find the dimensions in feet of his new garden
Answer by Theo(13342)

his garden is now a square with a perimeter of 60 feet.
this means that each side of the garden is equal to 60/4 = 15 feet.

he wants to turn it into a rectangular garden whose length is 3 feet shorter than its width.

if the width is going to stay the same as it was when the garden was square, then the width will be 15 feet.

2 * 15 feet = 30 feet and 3 feet shorter than that would be 27 feet.

the length of his garden will be 27 feet and the width will be 15 feet.

the perimeter of the new garden is 2 * 27 + 2 * 15 = 54 + 30 = 84 feet.